# Hölder curves with exotic tangent spaces

### v2

@@ 2.4. Tangents
Let $y,z\in r_{j}^{-1}(X-x)$ so that $\operatorname{dist}(y,Y)<d/3$ and $\operatorname{dist}(z,Z)<d/3$. Since $X$ is connected, there is another point $p\in r_{j}^{-1}(X-x)$ so that $p\in\overline{B}({\bf 0},R)$ and $\operatorname{dist}(p,Y)=d/3$; thus $\operatorname{dist}(p,Z)\geq 2d/3$. However, it follows that $\operatorname{dist}(p,T)\geq d/3$, which is a contradiction. ∎
Remark 2.6.
Note that since $\mathfrak{C}(\mathbb{R}^{N};{\bf 0})$ is sequentially compact, we have that if $X\in\mathfrak{C}(\mathbb{R}^{N})$ and if $x_{0}\in X$, then the sequence of sets $(n(X-x_{0}))_{n\in\mathbb{N}}$ has a subsequential limit in $\mathfrak{C}_{U}(\mathbb{R}^{N};{\bf 0})$. In particular, such a limit must be a tangent of $X$ at $x_{0}$, so the set of tangents to a nonempty closed set at a point contained in the set is always nonempty. More simply, every nonempty closed set has tangents at every point (and every tangent contains the origin $\bf 0$}).
@@ 2.5. Local cut-points
Let $p$ be a local cut-point of $X$ and let $r>0$ such that $C_{X}(p,r)\setminus\{p\}$ is disconnected. As a matter of notation, throughout this proof when we write $B_{X}(p,r)$ we mean $B(p,r)\cap X$, and similarly when we write $B_{Y}(f(p),r)$ we mean $B(f(p),r)\cap Y$. Let $\epsilon>0$ so that $\overline{B}_{Y}(f(p),\epsilon)\subset f(\overline{B}_{X}(p,r))$, and let $\delta\in(0,r)$ such that
@@
Let $A,B\subset C_{X}(p,r)\setminus\{p\}$ be a disjoint pair of nonempty closed subsets of $C_{X}(p,r)\setminus\{p\}$ so that $A\cup B=C_{X}(p,r)\setminus\{p\}$. Note that $A\cup\{p\}$ and $B\cup\{p\}$ are both closed sets in $X$. We claim that $A\cap C_{X}(p,\delta)\neq\emptyset$.
First, we show that $A\cup\{p\}$ is connected. If $A\cup\{p\}$ is not connected, then let $C,D$ be a disjoint pair of nonempty closed subsets of $A\cup\{p\}$ with $C\cup D=A\cup\{p\}$. Then since $A\cup\{p\}$ is closed in $X$, we have that $C$ and $D$ are both closed in $X$. Without loss of generality, assume that $p\in C$. Then the pair $B\cup C$ and $D$ is a disjoint pair of nonempty closed subsets of $C_{X}(p,r)$ with $(B\cup C)\cup D=C_{X}(p,r)$, which is a contradiction, so $A\cup\{p\}$ is connected.
@@
Furthermore, the sets on the right hand side form a disjoint pair of nonempty closed subsets of $C_{Y}(f(p),\epsilon)\setminus\{f(p)\}$ with $(f(A)\cap C_{Y}(f(p),\epsilon))\cup(f(B)\cap C_{Y}(f(p),\epsilon))=C_{Y}(f(p),\epsilon)\setminus\{f(p)\}$, so $f(p)$ is a local cut-point of $Y$. ∎
@@ 2.6. Self-similarity
If a function $f:\mathbb{R}^{N}\to\mathbb{R}^{M}$ is Lipschitz continuous, then we denote by $\operatorname{Lip}(f)$ (the Lipschitz norm of $f$) the smallest $L\geq 0$ so that for every pair of points $x,y\in\mathbb{R}^{N}$, $|f(x)-f(y)|\leq L|x-y|$. If $\operatorname{Lip}(f)<1$, then $f$ is a contraction. A map $f:\mathbb{R}^{N}\to\mathbb{R}^{M}$ is affine if there exists a linear map $L:\mathbb{R}^{N}\to\mathbb{R}^{M}$ so that for every $x\in\mathbb{R}^{N}$, $f(x)=L(x)+f({\bf 0})$. A map $f:\mathbb{R}^{N}\to\mathbb{R}^{N}$ is called a similarity if there exists $\lambda>0$ such that for every pair of points $x,y\in\mathbb{R}^{N}$, $|f(x)-f(y)|=\lambda|x-y|$. Every similarity is a composition of an orthogonal transformation, a scalar multiplication, and a translation; in particular similarity maps are affine. If $f:\mathbb{R}^{N}\to\mathbb{R}^{N}$ is a rotation-free and reflection-free similarity, then for every $x\in\mathbb{R}^{N}$, $f(x)=\operatorname{Lip}(f)x+f({\bf 0})$.
@@
A map $f:\mathbb{R}^{N}\to\mathbb{R}^{M}$ is affine if and only if there exists a linear transformation $L:\mathbb{R}^{N}\to\mathbb{R}^{M}$ so that for every $x,y\in\mathbb{R}^{N}$, f(x)f(y)=L(xy)f(x)-f(y)=L(x-y). In particular, if $f:\mathbb{R}^{N}\to\mathbb{R}^{N}$ is a rotation-free and reflection-free similarity, then for every $x,y\in\mathbb{R}^{N}$, $f(x)-f(y)=\operatorname{Lip}(f)(x-y)$.
@@
By a theorem of Schief [Sch94, Theorem 2.2], the OSC is equivalent to the strong open set condition (SOSC): if $K$ is the attractor of $\mathcal{F}$, then there exists a nonempty open set $U\subset\mathbb{R}^{N}$ for which the OSC holds so that $U\cap K\neq\emptyset$.
@@ 3. Tangents of Lipschitz curves at typical points
To estimate the second distance, fix $z\in\overline{C(x_{0},L,\alpha/2)}\cap\overline{B}(x_{0},\frac{3}{2}|y_{j}-x_{0}|)$. Since $y_{j}\notin C(x_{0},L,\alpha)$, we have $\operatorname{dist}(y_{j},L+x_{0})\geq\alpha|y_{j}-x_{0}|$. Moreover, $\operatorname{dist}(z,L+x_{0})\leq(\alpha/2)|z-x_{0}|$. Fix a point $p\in L+x_{0}$ so that $|z-p|\leq\tfrac{\alpha}{2}|z-x_{0}|$, so $|z-p|\leq\tfrac{3}{4}\alpha|y_{j}-x_{0}|$. Note that $|y_{j}-p|\geq\alpha|y_{j}-x_{0}|$. Thus, $|y_{j}-z|\geq\tfrac{\alpha}{4}|y_{j}-x_{0}|$. Hence, for each $j\in\mathbb{N}$,
@@
Let $f:[0,1]\to\mathbb{R}^{N}$ be Lipschitz and let $X=f([0,1])$. By [AO17, Thoerem 4.4], there exists $g:[0,1]\to X$, an essentially two-to-one Lipschitz parametrization with constant speed equal to $2\mathcal{H}^{1}(X)$. By Rademacher’s Theorem, for $\mathcal{H}^{1}$-a.e. $x\in X$, there exists $t\in g^{-1}(\{x\})$ so that $g^{\prime}(t)$ exists. Furthermore, by Lemma 3.2, $\mathcal{H}^{1}$-a.e. point $x\in X$ is flat in $X$. Let $x_{0}\in X$ be such that $x_{0}$ is flat in $X$ and there exists $t\in g^{-1}(\{x_{0}\})$ such that $g^{\prime}(t)$ exists. Let $T\in\operatorname{Tan}(X,x_{0})$, and let $L\subset\mathbb{R}^{N}$ be the line through the origin in the direction of $g^{\prime}(t)$.
First, we show that $T\subset L$. To this end, let $r_{j}\to 0$ be a sequence of positive scales so that $\frac{1}{r_{j}}(X-x_{0})\to T$ in the Attouch-Wets topology. Fix $\epsilon>0$ and $R>0$. By Lemma 3.2, there is $j_{0}\in\mathbb{N}$, such that for every $j\geq j_{0}$,
@@
Next, we show that $L\subset T$. Fix $\epsilon>0$ and $R>0$. There exists $h>0$ so that for every $s\in(-hR,hR)$,
@@
Moreover, there exists $j_{0}\in\mathbb{N}$ so that for every $j\geq j_{0}$ and every $s\in(-r_{j}R,r_{j}R)$ we get
@@ 4. A “statistically self-similar” carpet
Case 1. If $u(m)\neq u^{\prime}(m)$, then we have that $\phi_{u}^{\eta}((0,1)^{2})\subset\phi_{u(m)}^{\eta}((0,1)^{2})$ and $\phi_{u^{\prime}}^{\eta}((0,1)^{2})\subset\phi_{u^{\prime}(m)}^{\eta}((0,1)^{2})$ by the base case. Then, since $u(m)\neq u^{\prime}(m)$, we have by the induction hypothesis that $\phi_{u(m)}^{\eta}((0,1)^{2})\cap\phi_{u^{\prime}(m)}^{\eta}((0,1)^{2})=\emptyset$, so $\phi_{u}^{\eta}((0,1)^{2})\cap\phi_{u^{\prime}}^{\eta}((0,1)^{2})=\emptyset$.
Case 2. If $u(m)=u^{\prime}(m)$, then there exist distinct characters $i,j\in\mathcal{A}_{n}$ so that $u=u(m)j$ and $u^{\prime}=u^{\prime}(m)i$. By definition of $\phi_{u}^{\eta}$ we then have
@@
Let $\eta\in\mathscr{C}_{n}$, $w,v\in\mathcal{A}_{n}^{m}$, $x\in K_{w}^{\eta}$, and $y\in K_{v}^{\eta}$. If $|x-y|>2\sqrt{2}\cdot n^{-m}$, then $K_{w}^{\eta}\cap K^{\eta}_{v}=\emptyset$. If $|x-y|<n^{-m}$, then $K_{w}^{\eta}\cap K_{v}^{\eta}\neq\emptyset$.
@@
Thus, since $\partial[0,1]^{2}\subset X_{0}$, we have that $\partial[0,1]^{2}\subset X_{m}$ for every integer $m\geq 0$, and so since $K^{\eta}=\bigcap_{m=1}^{\infty}X_{m}$, we get $\partial[0,1]^{2}\subset K^{\eta}$. Similarly, for every $w\in\mathcal{A}_{n}^{*}$ we have that $\phi_{w}^{\eta}(\partial[0,1]^{2})\subset K_{w}^{\eta}$. Let $m\geq 1$ and let $v,w\in\mathcal{A}^{m}_{n}$. If $K_{v}^{\eta}\cap K_{w}^{\eta}\neq\emptyset$, then $\operatorname{diam}(K_{v}^{\eta}\cup K_{w}^{\eta})\leq 2\sqrt{2}n^{-m}$. Furthermore, if $K_{v}^{\eta}\cap K_{w}^{\eta}=\emptyset$, then $\phi_{v}^{\eta}(\partial[0,1]^{2})\cap\phi_{w}^{\eta}(\partial[0,1]^{2})=\emptyset$, so $\operatorname{dist}(\phi_{v}^{\eta}([0,1]^{2}),\phi_{w}^{\eta}([0,1]^{2}))\geq n^{-m}$, and the result follows as $K_{v}^{\eta}\subset\phi_{v}^{\eta}([0,1]^{2})$. ∎
@@ 5. Ahlfors regularity of carpets $K^{\eta}$
Note that $W(x,r)$ is a countable set since $\mathcal{A}_{n}^{*}$ is countable, and so $\bigcup_{w\in W(x,r)}K^{\eta}_{w}\in\Sigma$. Furthermore, for each point $y\in B(x,r)\cap K^{\eta}$ there is a word $\tau\in\mathcal{A}_{n}^{\mathbb{N}}$ with $\pi_{\eta}(\tau)=y$. Then since $B(x,r)\cap K^{\eta}$ is an open subset of $K^{\eta}$ and $\lim_{j\to\infty}\operatorname{diam}(K^{\eta}_{\tau(j)})=0,$ there is some $m\in\mathbb{N}$ large enough that $\tau(m)\in W(x,r)$. Since $y\in K^{\eta}_{\tau(m)}$, $\bigcup_{w\in W(x,r)}K^{\eta}_{w}=B(x,r)\cap K^{\eta}$, so we have that $\Sigma=\mathcal{B}(K^{\eta})$. ∎
Recall from §2.2 the $\sigma$-algebra $\Sigma_{n}$ on $\mathcal{A}^{\mathbb{N}}_{n}$ and the probability measure $\nu_{n}:\Sigma_{n}\to[0,1]$. We say a word $\tau\in\mathcal{A}_{n}^{\mathbb{N}}$ is an injective word if for every $\eta\in\mathscr{C}_{n}$, we have $\pi_{\eta}^{-1}(\{\pi_{\eta}(\tau)\})=\{\tau\}$. That is, $\tau$ is called an injective word if it always uniquely defines a point in $K^{\eta}$.
@@
Thus if a character $j\in\{4n-3,\dots,5n-6\}$ appears infinitely often in $\tau\in\mathcal{A}_{n}^{\mathbb{N}}$, then $\tau$ is injective; so the set of non-injective words is contained in the set of words for which every character $j\in\{4n-3,\dots,5n-6\}$ appears only finitely often. To see that this set is $\nu_{n}$-null, for each $m\in\mathbb{N}$ let $U_{m}:=\{\tau=i_{1}i_{2}\dots\in\mathcal{A}_{n}^{\mathbb{N}}:i_{m}>4n-4\}$. We have $\nu_{n}(U_{m})=\frac{n-2}{5n-6}$ for each $m$, and if $\{U_{m_{1}},\dots,U_{m_{j}}\}$ is a collection of these sets, then
@@
Fix $\eta\in\mathscr{C}_{n}$. We claim that for some $C_{n}>1$ the measure $\pi_{\eta}\#\nu_{n}$ defined on $\mathcal{B}(K^{\eta})$ by Lemma 5.2 satisfies
@@
Let $\{v_{1},\dots,v_{j}\}\subset\mathcal{A}_{n}^{m}$ be those words with $K_{v_{i}}^{\eta}\cap K_{\tau(m)}^{\eta}\neq\emptyset$. By Lemma 4.2, $j\leq 9$ and by Lemma 4.3, we have that $B(x,n^{-m})\cap K^{\eta}\subset\bigcup_{i=1}^{j}K^{\eta}_{v_{i}}$. We have also that $n^{-m}\geq r$, so
@@
6. Hölder parametrizations of carpets $K^{\eta}$

### v3

@@ 2.1. Tangents
Let $y,z\in r_{j}^{-1}(X-x)$ satisfy $\operatorname{dist}(y,Y)<d/3$ and $\operatorname{dist}(z,Z)<d/3$. Since $X$ is connected, there is another point $p\in r_{j}^{-1}(X-x)$ such that $p\in\overline{B}({\bf 0},R)$ and $\operatorname{dist}(p,Y)=d/3$; thus $\operatorname{dist}(p,Z)\geq 2d/3$. However, it follows that $\operatorname{dist}(p,T)\geq d/3$, which is a contradiction. ∎
2.2. Dendrites
Lemma 2.6.
Let $T_{1},T_{2}$ be two dendrites in $\mathbb{R}^{n}$ that intersect at a point. Then $T_{1}\cup T_{2}$ is a dendrite.
Proof.
2.3. Words
For each even integer $n\geq 4$ we denote
$\mathcal{A}_{n}=\{1,\dots,5n-6\}.$
2.4. Combinatorial graphs and trees
$E\subset\{\{v,v^{\prime}\}:v,v^{\prime}\in V\text{ and }v\neq v^{\prime}\}.$
$\text{Im}(G):=\bigcup_{\{v,v^{\prime}\}\in E}[v,v^{\prime}]$
A simple path in $G$ is a set $\gamma=\{\{v_{1},v_{2}\},\{v_{2},v_{3}\},\dots,\{v_{n-1},v_{n}\}\}\subset E$ such that for all distinct $i,j\in\{1,\dots,n\}$ we have $v_{i}\neq v_{j}$; in this case we say that $\gamma$ joins $v_{1}$, $v_{n}$. A graph $T=(V,E)$ is a combinatorial tree if for any distinct $v,v^{\prime}$ there exists a unique simple path $\gamma$ that joins $v$ with $v^{\prime}$.
@@ 2.5. Local cut-points
Let $p$ be a local cut-point of $X$ and let $r>0$ such that $C_{X}(p,r)\setminus\{p\}$ is disconnected. As a matter of notation, throughout this proof when we write $B_{X}(p,r)$ we mean $B(p,r)\cap X$, and similarly when we write $B_{Y}(f(p),r)$ we mean $B(f(p),r)\cap Y$. Let $\epsilon>0$ such that $\overline{B}_{Y}(f(p),\epsilon)\subset f(\overline{B}_{X}(p,r))$, and let $\delta\in(0,r)$ satisfy
@@
Let $A,B\subset C_{X}(p,r)\setminus\{p\}$ be a disjoint pair of nonempty closed subsets of $C_{X}(p,r)\setminus\{p\}$ such that $A\cup B=C_{X}(p,r)\setminus\{p\}$. Note that $A\cup\{p\}$ and $B\cup\{p\}$ are both closed sets in $X$. We claim that $A\cap C_{X}(p,\delta)\neq\emptyset$.
First, we show that $A\cup\{p\}$ is connected. If $A\cup\{p\}$ is not connected, then let $C,D$ be a disjoint pair of nonempty closed subsets of $A\cup\{p\}$ with $C\cup D=A\cup\{p\}$. Then since $A\cup\{p\}$ is closed in $X$, we have that $C$ and $D$ are both closed in $X$. Without loss of generality, assume that $p\in C$. Then the pair $B\cup C$ and $D$ is a disjoint pair of nonempty closed subsets of $C_{X}(p,r)$ with $(B\cup C)\cup D=C_{X}(p,r)$, which is a contradiction, hence $A\cup\{p\}$ is connected.
@@
Furthermore, the sets on the right hand side form a disjoint pair of nonempty closed subsets of $C_{Y}(f(p),\epsilon)\setminus\{f(p)\}$ with $(f(A)\cap C_{Y}(f(p),\epsilon))\cup(f(B)\cap C_{Y}(f(p),\epsilon))=C_{Y}(f(p),\epsilon)\setminus\{f(p)\}$, therefore $f(p)$ is a local cut-point of $Y$. ∎
@@ 2.6. Self-similarity
If a function $f:\mathbb{R}^{N}\to\mathbb{R}^{M}$ is Lipschitz continuous, then we denote by $\operatorname{Lip}(f)$ (the Lipschitz norm of $f$) the smallest $L\geq 0$ such that for every pair of points $x,y\in\mathbb{R}^{N}$, $|f(x)-f(y)|\leq L|x-y|$. If $\operatorname{Lip}(f)<1$, then $f$ is a contraction. A map $f:\mathbb{R}^{N}\to\mathbb{R}^{M}$ is affine if there exists a linear map $L:\mathbb{R}^{N}\to\mathbb{R}^{M}$ such that for every $x\in\mathbb{R}^{N}$, $f(x)=L(x)+f({\bf 0})$. A map $f:\mathbb{R}^{N}\to\mathbb{R}^{N}$ is called a similarity if there exists $\lambda>0$ such that for every pair of points $x,y\in\mathbb{R}^{N}$, $|f(x)-f(y)|=\lambda|x-y|$. Every similarity is a composition of an orthogonal transformation, a scalar multiplication, and a translation; in particular similarity maps are affine. If $f:\mathbb{R}^{N}\to\mathbb{R}^{N}$ is a rotation-free and reflection-free similarity, then for every $x\in\mathbb{R}^{N}$, $f(x)=\operatorname{Lip}(f)x+f({\bf 0})$.
@@
A map $f:\mathbb{R}^{N}\to\mathbb{R}^{M}$ is affine if and only if there exists a linear transformation $L:\mathbb{R}^{N}\to\mathbb{R}^{M}$ such that for every $x,y\in\mathbb{R}^{N}$, f(x)f(y)=L(xy)f(x)-f(y)=L(x-y). In particular, if $f:\mathbb{R}^{N}\to\mathbb{R}^{N}$ is a rotation-free and reflection-free similarity, then for every $x,y\in\mathbb{R}^{N}$, $f(x)-f(y)=\operatorname{Lip}(f)(x-y)$.
@@
By a theorem of Schief [Sch94, Theorem 2.2], the OSC is equivalent to the strong open set condition (SOSC): if $K$ is the attractor of $\mathcal{F}$, then there exists a nonempty open set $U\subset\mathbb{R}^{N}$ for which the OSC holds such that $U\cap K\neq\emptyset$.
@@
2.7. Self-similar sponges
(2.1) $\{S_{i}(y)=k^{-1}y+p_{i}:[0,1]^{N}\to[0,1]^{N}\}_{i=1}^{m},$
Lemma 2.9.
Proof.
$e:=S_{1}({\bf 0})-S_{2}({\bf 0})\in\{-k^{-1},0,k^{-1}\}^{N}.$
$S_{1}([0,1]^{N})\cap S_{2}([0,1]^{N})=S_{1}(I_{1}\times\dots\times I_{N}).$
For the case $l\geq 3$, we have that
@@ 3. Tangents of Lipschitz curves at typical points
To estimate the second distance, fix $z\in\overline{C(x_{0},L,\alpha/2)}\cap\overline{B}(x_{0},\frac{3}{2}|y_{j}-x_{0}|)$. Since $y_{j}\notin C(x_{0},L,\alpha)$, we have $\operatorname{dist}(y_{j},L+x_{0})\geq\alpha|y_{j}-x_{0}|$. Moreover, $\operatorname{dist}(z,L+x_{0})\leq(\alpha/2)|z-x_{0}|$. Fix a point $p\in L+x_{0}$ satisfying $|z-p|\leq\tfrac{\alpha}{2}|z-x_{0}|$, hence $|z-p|\leq\tfrac{3}{4}\alpha|y_{j}-x_{0}|$. Note that $|y_{j}-p|\geq\alpha|y_{j}-x_{0}|$. Therefore, $|y_{j}-z|\geq\tfrac{\alpha}{4}|y_{j}-x_{0}|$. Consequently, for each $j\in\mathbb{N}$,
@@
Let $f:[0,1]\to\mathbb{R}^{N}$ be Lipschitz and let $X=f([0,1])$. By [AO17, Thoerem 4.4], there exists $g:[0,1]\to X$, an essentially two-to-one Lipschitz parameterization with constant speed equal to $2\mathcal{H}^{1}(X)$. By Rademacher’s Theorem, for $\mathcal{H}^{1}$-a.e. $x\in X$, there exists $t\in g^{-1}(\{x\})$ such that $g^{\prime}(t)$ exists. Furthermore, by Lemma 3.2, $\mathcal{H}^{1}$-a.e. point $x\in X$ is flat in $X$. Let $x_{0}\in X$ be a point such that $x_{0}$ is flat in $X$ and there exists $t\in g^{-1}(\{x_{0}\})$ such that $g^{\prime}(t)$ exists. Let $T\in\operatorname{Tan}(X,x_{0})$, and let $L\subset\mathbb{R}^{N}$ be the line through the origin in the direction of $g^{\prime}(t)$.
First, we show that $T\subset L$. To this end, let $r_{j}\to 0$ be a sequence of positive scales satisfying $\frac{1}{r_{j}}(X-x_{0})\to T$ in the Attouch-Wets topology. Fix $\epsilon>0$ and $R>0$. By Lemma 3.2, there is $j_{0}\in\mathbb{N}$, such that for every $j\geq j_{0}$,
@@
Next, we show that $L\subset T$. Fix $\epsilon>0$ and $R>0$. There exists $h>0$ such that for every $s\in(-hR,hR)$,
@@
Moreover, there exists $j_{0}\in\mathbb{N}$ such that for every $j\geq j_{0}$ and every $s\in(-r_{j}R,r_{j}R)$ we have
@@ 4. A “statistically self-similar” carpet
Case 1. If $u(m)\neq u^{\prime}(m)$, then we have that $\phi_{u}^{\eta}((0,1)^{2})\subset\phi_{u(m)}^{\eta}((0,1)^{2})$ and $\phi_{u^{\prime}}^{\eta}((0,1)^{2})\subset\phi_{u^{\prime}(m)}^{\eta}((0,1)^{2})$ by the base case. Then, since $u(m)\neq u^{\prime}(m)$, we have by the induction hypothesis that $\phi_{u(m)}^{\eta}((0,1)^{2})\cap\phi_{u^{\prime}(m)}^{\eta}((0,1)^{2})=\emptyset$, therefore $\phi_{u}^{\eta}((0,1)^{2})\cap\phi_{u^{\prime}}^{\eta}((0,1)^{2})=\emptyset$.
Case 2. If $u(m)=u^{\prime}(m)$, then there exist distinct characters $i,j\in\mathcal{A}_{n}$ such that $u=u(m)j$ and $u^{\prime}=u^{\prime}(m)i$. By definition of $\phi_{u}^{\eta}$ we then have
@@
Let $\eta\in\mathscr{C}_{n}$, $w,v\in\mathcal{A}_{n}^{m}$, $x\in K_{w}^{\eta}$, and $y\in K_{v}^{\eta}$. If $|x-y|>2\sqrt{2}\cdot n^{-m}$, then $K_{w}^{\eta}\cap K^{\eta}_{v}=\emptyset$. If $|x-y|<n^{-m}$, then $K_{w}^{\eta}\cap K_{v}^{\eta}\neq\emptyset$. Furthermore, for each $\eta\in\mathscr{C}_{n}$ and every $w\in\mathcal{A}_{n}^{*}$, we have that $\phi^{\eta}_{w}(\partial[0,1]^{2})\subset K^{\eta}$. In particular, $\partial[0,1]^{2}\subset K^{\eta}$.
@@
Thus, since $\partial[0,1]^{2}\subset X_{0}$, we have that $\partial[0,1]^{2}\subset X_{m}$ for every integer $m\geq 0$, and since $K^{\eta}=\bigcap_{m=1}^{\infty}X_{m}$, we obtain $\partial[0,1]^{2}\subset K^{\eta}$. Similarly, for every $w\in\mathcal{A}_{n}^{*}$ we have that $\phi_{w}^{\eta}(\partial[0,1]^{2})\subset K_{w}^{\eta}$. Let $m\geq 1$ and let $v,w\in\mathcal{A}^{m}_{n}$. If $K_{v}^{\eta}\cap K_{w}^{\eta}\neq\emptyset$, then $\operatorname{diam}(K_{v}^{\eta}\cup K_{w}^{\eta})\leq 2\sqrt{2}n^{-m}$. Furthermore, if $K_{v}^{\eta}\cap K_{w}^{\eta}=\emptyset$, then $\phi_{v}^{\eta}(\partial[0,1]^{2})\cap\phi_{w}^{\eta}(\partial[0,1]^{2})=\emptyset$, thus $\operatorname{dist}(\phi_{v}^{\eta}([0,1]^{2}),\phi_{w}^{\eta}([0,1]^{2}))\geq n^{-m}$, and the result follows as $K_{v}^{\eta}\subset\phi_{v}^{\eta}([0,1]^{2})$. ∎
@@ 5. Ahlfors regularity of carpets $K^{\eta}$
Note that $W(x,r)$ is a countable set since $\mathcal{A}_{n}^{*}$ is countable, hence $\bigcup_{w\in W(x,r)}K^{\eta}_{w}\in\Sigma$. Furthermore, for each point $y\in B(x,r)\cap K^{\eta}$ there is a word $\tau\in\mathcal{A}_{n}^{\mathbb{N}}$ with $\pi_{\eta}(\tau)=y$. Then since $B(x,r)\cap K^{\eta}$ is an open subset of $K^{\eta}$ and $\lim_{j\to\infty}\operatorname{diam}(K^{\eta}_{\tau(j)})=0,$ there is some $m\in\mathbb{N}$ large enough that $\tau(m)\in W(x,r)$. Since $y\in K^{\eta}_{\tau(m)}$ and $\bigcup_{w\in W(x,r)}K^{\eta}_{w}=B(x,r)\cap K^{\eta}$, we have that $\Sigma=\mathcal{B}(K^{\eta})$. ∎
Recall from §2.3 the $\sigma$-algebra $\Sigma_{n}$ on $\mathcal{A}^{\mathbb{N}}_{n}$ and the probability measure $\nu_{n}:\Sigma_{n}\to[0,1]$. We say a word $\tau\in\mathcal{A}_{n}^{\mathbb{N}}$ is an injective word if for every $\eta\in\mathscr{C}_{n}$, we have $\pi_{\eta}^{-1}(\{\pi_{\eta}(\tau)\})=\{\tau\}$. That is, $\tau$ is called an injective word if it uniquely defines a point in $K^{\eta}$.
@@
Thus if a character $j\in\{4n-3,\dots,5n-6\}$ appears infinitely often in $\tau\in\mathcal{A}_{n}^{\mathbb{N}}$, then $\tau$ is injective; therefore the set of non-injective words is contained in the set of words for which every character $j\in\{4n-3,\dots,5n-6\}$ appears only finitely often. To see that this set is $\nu_{n}$-null, for each $m\in\mathbb{N}$ let $U_{m}:=\{\tau=i_{1}i_{2}\dots\in\mathcal{A}_{n}^{\mathbb{N}}:i_{m}>4n-4\}$. We have $\nu_{n}(U_{m})=\frac{n-2}{5n-6}$ for each $m$, and if $\{U_{m_{1}},\dots,U_{m_{j}}\}$ is a collection of these sets, then
@@
Fix $\eta\in\mathscr{C}_{n}$ and define the pushforward measure $\pi_{\eta}\#\nu_{n}:\mathcal{B}(K^{\eta})\to[0,1]$ by
$\pi_{\eta}\#\nu_{n}(A):=\nu_{n}(\pi_{\eta}^{-1}(A)).$
We claim that there exists $C_{n}>1$ such that
@@
Let $\{v_{1},\dots,v_{j}\}\subset\mathcal{A}_{n}^{m}$ be those words with $K_{v_{i}}^{\eta}\cap K_{\tau(m)}^{\eta}\neq\emptyset$. By Lemma 4.2, $j\leq 9$ and by Lemma 4.3, we have that $B(x,n^{-m})\cap K^{\eta}\subset\bigcup_{i=1}^{j}K^{\eta}_{v_{i}}$. We have also that $n^{-m}\geq r$, hence
@@
6. Hölder parameterizations of carpets $K^{\eta}$